If the​ cost, C​, for manufacturing x units of a certain product is given by C=x^2-20x+55, find the number of units manufactured at a cost of $9955.

Answer by Boreal(15235)

9955=x^2-20x+55
x^2-20x-9900=0
(x-110)(x+90)=0
only positive root is x=110 units ANSWER
12100-2200+55=9955.